# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ Abstract
Dynamic Text-Attributed Graphs (DyTAGs), which intricately integrate structural, temporal, and textual attributes, are crucial for modeling complex real-world systems. However, most of the existing DyTAG datasets exhibit poor textual quality, which severely limits their utility for DyTAG generation tasks requiring semantically rich inputs. Additionally, prior work mainly focuses on discriminative tasks on DyTAGs, resulting in a lack of standardized task formulations and evaluation protocols tailored for DyTAG generation. To address these critical issues, we propose Generative DyTAG Benchmark (GDGB), which comprises eight meticulously curated DyTAG datasets with high-quality textual features for both nodes and edges, overcoming limitations of prior datasets. Building on GDGB, we define two novel DyTAG generation tasks: Transductive Dynamic Graph Generation (TDGG) and Inductive Dynamic Graph Generation (IDGG). TDGG transductively generates a target DyTAG based on the given source and destination node sets, while the more challenging IDGG introduces new node generation to inductively model the dynamic expansion of real-world graph data. To enable holistic evaluation, we design multifaceted metrics that assess the structural, temporal, and textual quality of the generated DyTAGs. We further propose GAG-General, an LLM-based multi-agent generative framework tailored for reproducible and robust benchmarking of DyTAG generation. Experimental results demonstrate that GDGB enables rigorous evaluation of TDGG and IDGG, with key insights revealing the critical interplay of structural and textual features in DyTAG generation. These findings establish GDGB as a foundational resource for advancing generative DyTAG research and unlocking further practical applications in DyTAG generation. GDGB datasets, source codes, and leaderboards are available at here.
@@ 1 Introduction
Dynamic graph-structured data is ubiquitous, spanning various domains, such as social networks huang2022ttergm ; sun2022aligning , recommendation systems zhang2022dynamic ; tang2023dynamic , and citation networks skarding2021foundations ; geng2022modeling . Notably, real-world dynamic evolution scenarios inherently contain rich (textual) features. For example, user posts and interactions such as comments and reposts on social networks, and consumer reviews of products on e-commerce platforms. This motivates the construction of dynamic text-attributed graphs (DyTAGs), which systematically integrate structural/temporal dynamics and textual attributes DTGB .
@@
Existing representative DyTAG benchmarks, such as DTGB DTGB , demonstrate that state-of-the-art Dynamic Graph Neural Networks (DGNNs) achieve significant performance gains in discriminative tasks (e.g., link prediction, node retrieval, and edge classification) by leveraging DyTAGs’ textual features. However, despite the growing interest in generative tasks within computer vision (CV) wang2021generative ; raut2024generative and natural language processing (NLP) xu2024large ; blease2024generative , as well as the increasing demand for graph generation in domains such as drug discovery luo20213d ; martinelli2022generative , graph augmentation ding2022data ; meng2023generative , and privacy-preserving frameworks miao2023graph ; hetemporal2025 , DyTAG generation remains underexplored. For instance, the DTGB demonstrates shallow exploration of DyTAG generative tasks, focusing narrowly on simplistic textual relation generation while neglecting the structural and temporal dynamics of graph evolution DTGB . This gap underscores the increasing criticality of generating high-quality DyTAGs.
@@
The primary obstacle hindering progress in DyTAG generation lies in the absence of high-quality, text-rich DyTAG benchmarks. Current datasets and benchmarks face two critical limitations: 1) Lack of high-quality (textual) attributes: Traditional dynamic graph datasets fundamentally lack node/edge features, relying solely on topological and temporal information DGB ; TGB-seq . This scarcity of attributes poses substantial challenges for generative models that require rich feature inputs. Meanwhile, existing DyTAG datasets from DTGB exhibit poor textual quality, with (source) node texts typically limited to usernames or email addresses, lacking semantic richness DTGB . 2) Lack of standardized DyTAG generative task formulations and evaluation protocols: Current dynamic graph generative models mainly rely on structural and temporal information TIGGER ; TG-GAN ; G2A2 , necessitating the development of new task formulations and holistic evaluation metrics tailored to DyTAG’s textual characteristics. Furthermore, most approaches prioritize direct generation of the final target graph Dymond ; VRDAG , diverging from the incremental and expansive growth patterns observed in real-world dynamic graph generation scenarios GAG . To address these issues, we propose Generative DyTAG Benchmark (GDGB) from three aspects: dataset construction, task and metric definition, and framework design.
@@
Framework. Given the text-rich nature of DyTAGs, we propose GAG-General, an LLM-based multi-agent framework tailored for DyTAG generation tasks, extending the implementation of GAG GAG . Unlike the bipartite-graph-specific GAG, GAG-General demonstrates superior generalization across diverse DyTAG scenarios, without requiring domain-specific customization. Furthermore, we implement both TDGG and IDGG tasks within GAG-General, integrating the proposed holistic evaluation metrics to ensure reproducible DyTAG generation and robust benchmarking.
@@
  • First Generative DyTAG Benchmark. To the best of our knowledge, GDGB is the first generative DyTAG benchmark, including eight high-quality datasets for DyTAG generation.
@@ 2 Related Work
Discriminative Dynamic Graph Learning. Recent years have witnessed significant advances in dynamic graph learning over discrete-time dynamic graphs (DTDGs) and continuous-time dynamic graphs (CTDGs). Substantial progress has been made through various DGNNs EvolveGCN ; you2022roland ; TGN ; DyGLib ; peng2025tidformer . These models demonstrate strong performance in discriminative tasks, including dynamic link prediction, node retrieval, etc. Concurrently, the community develops diverse benchmarks to standardize evaluation DGB ; DyGLib ; TGB ; TGB-seq . Notable initiatives include DyGLib DyGLib benchmarking on classical dynamic graph datasets, TGB TGB introducing large-scale graphs for scalability challenges, and TGB-Seq TGB-seq focusing on complex sequential patterns. However, current benchmarks primarily provide basic structural-temporal information, with limited datasets containing simplistic statistical attributes for nodes or edges DTGB . To address this, DTGB DTGB proposes DyTAG datasets and incorporates textual features through BERT embeddings BERT , showing measurable performance gains across multiple discriminative tasks DTGB . Nevertheless, the information bottleneck caused by transformed BERT embeddings inevitably compromises original semantic richness, leaving open the challenge of effectively utilizing textual semantics for enhanced DyTAG learning.
@@
Generative Dynamic Graph Learning. Compared to its discriminative counterpart, generative dynamic graph learning is still in its nascent stages. Early approaches primarily focus on DTDGs, including Dymond ; taggen ; G2A2 ; D2G2 . Recent advances have shifted toward CTDG generation, enabling more refined temporal modeling that better aligns with real-world application needs DGGEN ; TG-GAN ; TIGGER . Contemporary approaches begin exploring feature-supportive generation: VRDAG VRDAG implements node attribute generation through graph-based variational autoencoders (VAEs), while DG-Gen DGGEN models edge attributes via joint conditional probability distributions. More recently, GAG GAG , an LLM-based multi-agent framework, enables simulation-based generation for social bipartite DyTAGs. However, this field significantly lacks standardized benchmarks for comprehensive and effective evaluation of dynamic graph generation tasks. First, in terms of task design, existing methods prioritize direct generation of the final target graph Dymond ; taggen ; VRDAG , which deviates from the expansive growth patterns observed in real-world dynamic graph generation scenarios GAG . Second, the definition of evaluation metrics lacks multidimensional, holistic assessments of graph structure, temporal dynamics, and attribute coherence taggen ; VRDAG ; GAG . Therefore, when extending generative tasks to DyTAG generation, establishing a robust benchmark tailored to DyTAG’s dynamic and text-rich characteristics becomes a critical research need.
@@ 3 Proposed Datasets
Limitations of Existing Datasets. Existing dynamic graph datasets, while valuable for structural and temporal modeling, suffer from critical shortcomings regarding textual attributes that hinder high-quality DyTAG generation. DTGB DTGB makes the first step where both information is included, however, as illustrated in Figure 2, the average node and edge text lengths in DTGB are notably shorter and semantically trivial. For instance, six out of eight datasets of DTGB contain (source) node texts that are typically just identifiers like email addresses or usernames (Figure 2 Left). While the two datasets from the Stack platform attempt to enrich texts by recording user locations and introductions, half of the user node fields remain empty or nonsensical. Furthermore, edge texts in datasets like GDELT and ICEWS1819 are overly brief (Figure 2 Right), offering insufficient contextual information for text generation. These limitations of existing datasets hinder the advancement of DyTAG generation research. A comprehensive comparison of more existing datasets is provided in Section A.4.
Our Proposed High-Quality DyTAG Datasets. To overcome the aforementioned textual quality issues and enable DyTAG generation tasks, we propose Generative DyTAG Benchmark, GDGB, consisting of eight carefully curated DyTAG datasets. These datasets are sourced from diverse real-world domains, including e-commerce, social networks, biography networks, etc., encompassing both bipartite (four datasets) and non-bipartite (four datasets) graphs. Detailed statistics are presented in Table 1. A key focus during the collection and preprocessing of these datasets is to ensure that all source/destination nodes and interaction edges possess rich, semantic textual attributes. For example, Sephora includes: 1) user node texts detailing user appearances and review histories, 2) product node texts describing product brands, ingredients, and ratings, and 3) edge texts containing user detailed reviews. The textual completeness of GDGB provides a robust foundation for developing DyTAG generation models capable of producing both realistic structures and coherent text.
Table 1: Statistics of proposed DyTAG datasets of GDGB. See Section A.1 for more details.
@@
WikiRevision 75,622/3,204 2,778,732 2 276,6153 Web Interaction Node & Edge
@@
Figure 1: Left: Average node text lengths on GDGB datasets and DTGB datasets. In non-bipartite graphs, text lengths are averaged across all nodes. For bipartite graphs, averages are calculated for source nodes. Right: Average edge text lengths on our datasets and DTGB datasets.
Figure 2: Left: Comparison of Degree MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG. Right: Comparison of Spectra MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG. (The dataset order is consistent with Figure 2).
Generative Task Performance Validation. To further validate the impact of textual quality on dynamic graph generation tasks, we conduct experiments using the latest models that support dynamic graph generation with node/edge features. We select two representative dynamic graph generation models: VRDAG VRDAG , which supports node features, and DG-Gen DGGEN , which supports edge features. Using public implementations from the original works, we evaluate the structural quality of generated graphs under both conditions (w/ and w/o textual features). Following DTGB’s setting DTGB , we employ the BERT-base-uncased model BERT to encode raw texts into initial representations and apply the maximum mean discrepancy (MMD) MMD metric to quantify differences in degree/spectra distribution (Degree/Spectra MMD) between generated and ground-truth graphs. Ideally, if textual features are helpful, the generated graphs with these features would have higher quality with lower MMD. Results in Figure 2 and 7 show that on our proposed GDGB, incorporating node or edge texts significantly reduces Degree/Spectra MMD, indicating improved graph structural quality. Conversely, in DTGB, textual features even negatively impact performance in half of the cases—a phenomenon more pronounced in the node-feature-specific VRDAG model for the poor quality of node texts. This further proves the inherent textual quality issues in DTGB.
@@ 4.1 Generative Task Design
Transductive Dynamic Graph Generation. Starting with an initial seed DyTAG $\mathcal{G}_{0}$, TDGG generates a final DyTAG $\mathcal{G}_{K}$ based on the given source and destination node sets, maintaining the transductive assumption that all nodes are known as prior knowledge. Thus, the goal of TDGG is to conduct destination node selection and edge generation, which naturally integrates traditional discriminative tasks (e.g., node retrieval and edge classification) in a generative paradigm. The generated DyTAG $\mathcal{G}_{K}$ in TDGG is expected to maintain structural, temporal, and textual similarity to the ground-truth DyTAG, reflecting high-quality generation. The illustration of TDGG is shown in Figure 3 (b), demonstrating that the user will review the powder at $t^{\prime}$, along with a generated negative review and a rating of 1.
Inductive Dynamic Graph Generation. Different from TDGG, the more challenging IDGG extends the transductive setting by introducing inductive modeling of new node generation during graph evolution. Starting with an initial seed DyTAG $\mathcal{G}_{0}$, IDGG generates a final DyTAG $\mathcal{G}_{K}$ based on the dynamically evolving source and destination node sets with newly generated nodes. Due to the incorporation of inductive node generation, the generated DyTAG $\mathcal{G}_{K}$ in IDGG is expected to preserve the structural properties of the ground-truth DyTAG, while ensuring that all newly added nodes and edges contain high-quality, semantically coherent textual attributes, thereby successfully modeling the dynamic expansion of real-world graph data. The illustration of IDGG is shown in Figure 3 (c), showcasing that the newly generated user will review the newly generated lipstick at $t^{\prime\prime}$ with a generated positive review and a rating of 4.

### v2

@@ Abstract
Dynamic Text-Attributed Graphs (DyTAGs), which intricately integrate structural, temporal, and textual attributes, are crucial for modeling complex real-world systems. However, most existing DyTAG datasets exhibit poor textual quality, which severely limits their utility for generative DyTAG tasks requiring semantically rich inputs. Additionally, prior work mainly focuses on discriminative tasks on DyTAGs, resulting in a lack of standardized task formulations and evaluation protocols tailored for DyTAG generation. To address these critical issues, we propose Generative DyTAG Benchmark (GDGB), which comprises eight meticulously curated DyTAG datasets with high-quality textual features for both nodes and edges, overcoming limitations of prior datasets. Building on GDGB, we define two novel DyTAG generation tasks: Transductive Dynamic Graph Generation (TDGG) and Inductive Dynamic Graph Generation (IDGG). TDGG transductively generates a target DyTAG based on the given source and destination node sets, while the more challenging IDGG introduces new node generation to inductively model the dynamic expansion of real-world graph data. To enable holistic evaluation, we design multifaceted metrics that assess the structural, temporal, and textual quality of the generated DyTAGs. We further propose GAG-General, an LLM-based multi-agent generative framework tailored for reproducible and robust benchmarking of DyTAG generation. Experimental results demonstrate that GDGB enables rigorous evaluation of TDGG and IDGG, with key insights revealing the critical interplay of structural and textual features in DyTAG generation. These findings establish GDGB as a foundational resource for advancing generative DyTAG research and unlocking further practical applications in DyTAG generation. The dataset and source code are available at https://github.com/Lucas-PJ/GDGB-ALGO.
@@ 1 Introduction
Dynamic graph-structured data is ubiquitous, spanning various domains, such as social networks (Huang et al., 2022; Sun et al., 2022), recommendation systems (Zhang et al., 2022; Tang et al., 2023), and citation networks (Skarding et al., 2021; Geng et al., 2022). Notably, real-world dynamic evolution scenarios inherently contain rich (textual) features. For example, user posts and interactions such as comments and reposts on social networks, and consumer reviews of products on e-commerce platforms. This motivates the construction of dynamic text-attributed graphs (DyTAGs), which systematically integrate structural/temporal dynamics and textual attributes (Zhang et al., 2024a).
@@
Existing representative DyTAG benchmarks, such as DTGB (Zhang et al., 2024a), demonstrate that state-of-the-art Dynamic Graph Neural Networks (DGNNs) achieve significant performance gains in discriminative tasks (e.g., link prediction, node retrieval, and edge classification) by leveraging DyTAGs’ textual features. Growing interest in generative tasks spans computer vision (CV) (Wang et al., 2021b; Raut and Singh, 2024) and natural language processing (NLP) (Xu et al., 2024; Blease et al., 2024), with increasing demand for graph generation in domains such as drug discovery (Luo et al., 2021; Martinelli, 2022), graph augmentation (Ding et al., 2022; Meng et al., 2023), and privacy-preserving frameworks (Miao et al., 2023; He et al., 2025). However, DyTAG generation remains underexplored. For instance, DTGB focuses narrowly on simplistic textual relation generation while neglecting the rich interplay of textual, structural, and temporal dynamics in graph evolution (Zhang et al., 2024a). This gap underscores the increasing criticality of generating high-quality DyTAGs.
@@
The primary obstacle hindering progress in DyTAG generation lies in the absence of high-quality, text-rich DyTAG benchmarks. Current datasets and benchmarks face two critical limitations: 1) Lack of high-quality (textual) attributes: Traditional dynamic graph datasets fundamentally lack node/edge features, relying solely on topological and temporal information (Poursafaei et al., 2022b). This scarcity of attributes poses substantial challenges for generative models that require rich feature inputs. Meanwhile, existing DyTAG datasets from DTGB exhibit poor textual quality, with (source) node texts typically limited to usernames or email addresses, lacking semantic richness (Zhang et al., 2024a), which severely restricts the development of DyTAG learning. 2) Lack of standardized DyTAG generative task formulations and evaluation protocols: Current dynamic graph generative models mainly rely on structural and temporal information (Gupta et al., 2022; Dey et al., 2024), necessitating the development of new task formulations and holistic evaluation metrics tailored to DyTAG’s textual characteristics. Furthermore, most approaches prioritize direct generation of the final target graph (Zeno et al., 2021; Li et al., 2024), diverging from the incremental and expansive growth patterns observed in real-world dynamic graph generation scenarios (Ji et al., 2025). Hence, to address these issues, we propose the Generative DyTAG Benchmark (GDGB) from the following three aspects: 1) dataset construction, 2) task & metric definition, and 3) generative framework design.
@@
Framework. Given the text-rich nature of DyTAGs, we propose GAG-General, an LLM-based multi-agent framework tailored for DyTAG generation tasks, extending the implementation of GAG (Ji et al., 2025). Unlike the bipartite-graph-specific GAG, GAG-General demonstrates superior generalization across diverse DyTAG scenarios, without requiring domain-specific customization. Furthermore, we implement both TDGG and IDGG tasks within GAG-General, integrating the proposed holistic evaluation metrics to ensure reproducible DyTAG generation and robust benchmarking.
@@
  • First Generative DyTAG Benchmark. To the best of our knowledge, GDGB is the first generative DyTAG benchmark, including eight high-quality datasets tailored for DyTAG generation.
@@ 2 Related Work
Discriminative Dynamic Graph Learning. Recent years have witnessed significant advances in dynamic graph learning over discrete-time dynamic graphs (DTDGs) and continuous-time dynamic graphs (CTDGs). Substantial progress has been made through various DGNNs (Rossi et al., 2020; Peng et al., 2025). These models demonstrate strong performance in discriminative tasks, including dynamic link prediction, node retrieval, etc. Concurrently, the community develops diverse benchmarks to standardize evaluation (Poursafaei et al., 2022b; Huang et al., 2024). Notable initiatives include DyGLib (Yu et al., 2023) benchmarking on classical dynamic graph datasets, TGB (Huang et al., 2024) introducing large-scale graphs for scalability challenges, and TGB-Seq (Yi et al., 2025) focusing on complex sequential patterns. However, current benchmarks primarily provide basic structural-temporal information, with limited datasets containing simplistic statistical attributes for nodes or edges (Zhang et al., 2024a). To address this, DTGB (Zhang et al., 2024a) proposes DyTAG datasets and incorporates textual features through BERT embeddings (Devlin et al., 2019), showing measurable performance gains across multiple discriminative tasks (Zhang et al., 2024a). Nevertheless, the information bottleneck caused by transformed BERT embeddings inevitably compromises original semantic richness, leaving open the challenge of effectively utilizing textual semantics for enhanced DyTAG learning.
@@
Generative Dynamic Graph Learning. Compared to its discriminative counterpart, generative dynamic graph learning is still in its nascent stages. Early approaches primarily focus on DTDGs, including (Zeno et al., 2021; Zhou et al., 2020; Dey et al., 2024). Recent advances have shifted toward CTDG generation, enabling more refined temporal modeling that better aligns with real-world application needs (Zhang et al., 2021; Gupta et al., 2022). Contemporary approaches begin exploring feature-supportive generation: VRDAG (Li et al., 2024) implements node attribute generation through graph-based variational autoencoders (VAEs), while DG-Gen (Hosseini et al., 2024) models edge attributes via joint conditional probability distributions. More recently, GAG (Ji et al., 2025), an LLM-based multi-agent framework, enables simulation-based generation for social bipartite DyTAGs. However, this field significantly lacks standardized benchmarks for comprehensive and effective evaluation of dynamic graph generation tasks. First, in terms of task design, existing methods prioritize direct generation of the final target graph (Zeno et al., 2021; Zhou et al., 2020; Li et al., 2024), which deviates from the expansive growth patterns observed in real-world dynamic graph generation scenarios (Ji et al., 2025). Second, the definition of evaluation metrics lacks multidimensional, holistic assessments of graph structure, temporal dynamics, and attribute coherence (Zhou et al., 2020; Li et al., 2024; Ji et al., 2025). Therefore, when extending generative tasks to DyTAG generation, establishing a robust benchmark tailored for DyTAG’s dynamic and text-rich characteristics becomes a critical need.
@@ 3 Proposed Datasets
Table 1: Statistics of proposed DyTAG datasets of GDGB. See Section˜C.1 for more details.
@@
WikiRevision 75,622/3,204 2,778,732 2 2,766,153 Web Interaction Node & Edge
@@
Figure 1: Comparison of node and edge texts across GDGB and DTGB datasets in terms of length, perplexity (PPL (Reversed)), and LLM-based rating.
Limitations of Existing Datasets. Existing dynamic graph datasets, though valuable for structural and temporal modeling, suffer from critical limitations in textual attributes that hinder high-quality DyTAG generation. DTGB (Zhang et al., 2024a) represents the first effort to incorporate both node and edge texts. However, as shown in Figure˜1, the node and edge texts in DTGB are notably short and semantically shallow in terms of length, perplexity (PPL), and LLM-based rating. GDGB achieves significant quality advantages over DTGB in five out of six dimensions (see detailed results for each dimension in Section˜C.5). For instance, in six of DTGB’s eight datasets, (source) node texts are typically mere identifiers such as email addresses or usernames (Figure˜5, left). While the two Stack-platform datasets attempt to enrich node texts with user locations and bios, half of these fields remain empty or meaningless. Moreover, edge texts in datasets like GDELT and ICEWS1819 are excessively brief (Figure˜5, right), providing insufficient context for text generation. Despite DTGB having longer average edge text length, the semantic quality of DTGB’s texts—as indicated by PPL and LLM-based rating—is significantly lower than GDGB’s. These limitations of current datasets impede progress in DyTAG generation research. A comprehensive comparison of additional existing datasets is provided in Section˜C.4.
Our Proposed High-Quality DyTAG Datasets. To overcome the aforementioned textual quality issues and enable DyTAG generation tasks, we propose Generative DyTAG Benchmark, GDGB, consisting of eight carefully curated DyTAG datasets. These datasets are sourced from diverse real-world domains, including e-commerce, social networks, biography networks, etc., encompassing both bipartite (4 datasets) and non-bipartite (4 datasets) graphs. Detailed statistics are presented in Table˜1. A key focus during the collection and preprocessing of these datasets is to ensure that all source/destination nodes and interaction edges possess rich, semantic textual attributes. For example, Sephora includes: 1) user node texts detailing user appearances and review histories, 2) product node texts describing product brands, ingredients, and ratings, and 3) edge texts containing user detailed reviews. The textual completeness of GDGB provides a robust foundation for developing DyTAG generation models capable of producing both realistic structures and coherent text.
Generative Task Performance Validation. To assess the impact of textual quality on dynamic graph generation, we evaluate two latest feature-supportive models—VRDAG (Li et al., 2024) (node features) and DG-Gen (Hosseini et al., 2024) (edge features)—using their public implementations. Following DTGB’s setup (Zhang et al., 2024a), we encode raw texts with BERT-base-uncased (Devlin et al., 2019) and measure structural fidelity via Degree/Spectra MMD (Xiang et al., 2022) to quantify distributional differences between generated and ground-truth graphs. Lower MMD indicates better structural quality when textual features are effective. Results in Figures˜7 and 7 show that, on GDGB, incorporating node or edge texts significantly reduces MMD, improving structural similarity. In contrast, on DTGB, textual features degrade performance in half the cases—especially for VRDAG—due to poor text quality. This highlights inherent limitations in DTGB. Beyond these experiments, we further validate GDGB’s advantages through applying our generative framework (proposed in Section˜4.3) to downstream generation tasks, providing a comprehensive demonstration of the advantages and distinctions of GDGB over DTGB in Section˜C.5.
@@ 4.1 Generative Task Design
Transductive Dynamic Graph Generation. Starting with an initial seed DyTAG $\mathcal{G}_{0}$, TDGG generates a final DyTAG $\mathcal{G}_{K}$ based on the given source and destination node sets, maintaining the transductive assumption that all nodes are known as prior knowledge. Thus, the goal of TDGG is to conduct destination node selection and edge generation, which naturally integrates traditional discriminative tasks (e.g., node retrieval and edge classification) in a generative paradigm. The generated DyTAG $\mathcal{G}_{K}$ in TDGG is expected to maintain structural, temporal, and textual similarity to the ground-truth DyTAG, reflecting high-quality generation. The illustration of TDGG is shown in Figure˜2 (b), demonstrating that the user will review the powder at $t^{\prime}$, along with a generated negative review and a rating of 1.
Inductive Dynamic Graph Generation. Different from TDGG, the more challenging IDGG extends the transductive setting by introducing inductive modeling of new node generation during graph evolution. Starting with an initial seed DyTAG $\mathcal{G}_{0}$, IDGG generates a final DyTAG $\mathcal{G}_{K}$ based on the dynamically evolving source and destination node sets with newly generated nodes. Due to the incorporation of inductive node generation, the generated DyTAG $\mathcal{G}_{K}$ in IDGG is expected to preserve the structural properties of the ground-truth DyTAG, while ensuring that all newly added nodes and edges contain high-quality, semantically coherent textual attributes, thereby successfully modeling the dynamic expansion of real-world graph data. The illustration of IDGG is shown in Figure˜2 (c), showcasing that the newly generated user will review the newly generated lipstick at $t^{\prime\prime}$ with a generated positive review and a rating of 4.